OpenStudy (anonymous):

Ann and Tom want to establish a fund for their grandsons college education. What lump sum must they deposit at a 9.7% annual interest rate, compounded quarterly, to have 50,000 in the fund at the end of 10 years. Ok, so i did the math kinda... but got a really off number of 1.2707..... not sure where I went wrong

OpenStudy (wolf1728):

The formula is here: http://www.1728.org/compint3.htm Principal = Total / (1 + rate/n)^n*years where n = number of compounding periods per year which in this case is 4 Principal = 50,000 / (1 + (.097/4))^4*10 Principal = 50,000 / (1.02425)^4*10 Principal = 50,000 / 2.6075875871 Principal = 19,174.81 There's a LOT of places to make mistakes don't you think?

OpenStudy (anonymous):

Wow, well thank you! haha I see where I mistyped it in my calculator

OpenStudy (wolf1728):

It's okay - glad I could help you out.